# Magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances in hot and dense strange hadronic matter

### v1

@@ IV.2 In-medium magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances
A study based on QMC as well as MQMC, for octet baryons, shows that we can expect sizable influence on effective baryonic moments due to temperature above $T=150$ MeV at lower medium densities mqmc . Speaking of correlations, the authors in ref. seventyfour have discussed a strong correlation of the baryonic magnetic moment with that of the bag radius of a baryon, as the two quantities have a similar behavior as density dependent functions. Since this is the first work on medium modifications of $\frac{1}{2}^{-}$ baryon resonances, so referring to previous works on in-medium density effects on magnetic moments of baryons using the current formalism sixtytwo ; fiftyfour and QMC tsushima2 , might facilitate the purpose of a comparative study for future work on the magnetic moments of $\frac{1}{2}^{-}$ baryon resonances.
V Summary
@@
To summarize, in the present work we have investigated the modifications induced on the masses as well as the magnetic moments of $J_{p}={\frac{1}{2}}^{-}$ baryon resonances in isospin asymmetric strange hadronic matter. In light of chiral $SU(3)$ quark mean field model the in-medium effects are introduced and we computed the effective baryon masses by using the masses of constituent quarks. The mass adjusted quark magnetic moments are then calculated, to include relativistic quark confinement, by using the baryon masses as inputs. In hadronic medium the in-medium masses of lights quarks, $u$ $d$, vary significantly in comparison to the observed changes in medium modified masses of $s$ quark. In the $\chi$CQM formalism, we then computed the baryonic magnetic moments from the contributions due to valence quarks, sea quarks and orbital angular momentum due of sea quarks. As derived from the results is that, the medium modifications are more prominent for the baryons comprised of light $u$ or $d$ quarks, than for hyperons. For strange matter, the in-medium constituent quark masses, effective baryonic masses as well as the effective baryonic magnetic moments suffer appreciable enhancement in their in-medium behaviour. By carrying out this study at finite temperatures and finite $i_{a}$ values of the hadronic medium, we have noted a fine shift in the in-medium baryonic magnetic moments, although the temperature effects have been negligible below $T=150$ MeV.
@@
Table 3: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of symmetric nuclear matter ($i_{a}=0$ and $f_{s}=0$) at a medium temperature $T=0$ MeV.
@@
${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.280 -0.389 0.213 2.104 2.792 -0.475 0.032 2.349
${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -1.459 0.263 -0.185 -1.382 -1.676 0.303 -0.003 -1.376
@@
Table 4: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of asymmetric nuclear matter ($i_{a}=0.5$ and $f_{s}=0$) at a medium temperature $T=0$ MeV.
@@
Table 5: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of symmetric hyperonic matter ($i_{a}=0$ and $f_{s}=0.3$) at a medium temperature $T=0$ MeV.
@@
Table 6: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of asymmetric hyperonic matter ($i_{a}=0.5$ and $f_{s}=0.3$) at a medium temperature $T=0$ MeV.
@@
$\mu_{\text{val}}$ $\mu_{\text{sea}}$ $\mu_{\text{orbit}}$ $\mu_{B}$ $\mu_{B}$ (GBE) $\mu_{B}$ (OGE) $\mu_{B}$ (GBE) $\mu_{B}$ $\mu_{B}$ $\mu_{B}$
@@
VI APPENDIX
By minimizing the thermopotential in eq. 27, we derive the equations of motion for scalar fields $\sigma$, $\zeta$, the dilaton field $\chi$, scalar iso-vector field $\delta$ and the vector fields $\omega,\rho,\phi$ in view of eq. 29 as in fiftyfour
[…]
$\,+\,\frac{\chi^{2}}{\chi_{0}^{2}}m_{\pi}^{2}f_{\pi}-\left(\frac{\chi}{\chi_{0}}\right)^{2}m_{\omega}\omega^{2}\frac{\partial m_{\omega}}{\partial\sigma}\,-\left(\frac{\chi}{\chi_{0}}\right)^{2}m_{\rho}\rho^{2}\frac{\partial m_{\rho}}{\partial\sigma}\,+\,\sum_{i}\frac{\partial M_{i}^{\ast}}{\partial\sigma}\rho_{i}^{s}=0,~{}~{}~{}~{}~{}$ (53)
[…]
$+\,\frac{\chi^{2}}{\chi_{0}^{2}}\left(\sqrt{2}m_{K}^{2}f_{K}\,-\,\frac{1}{\sqrt{2}}m_{\pi}^{2}f_{\pi}\right)+\,\sum_{i=\Lambda,\Sigma^{\pm,0},\Xi^{-,0}}\frac{\partial M_{i}^{\ast}}{\partial\zeta}\rho_{i}^{s}=0,~{}~{}~{}~{}~{}=0,$ (54)
$\frac{\partial\Omega}{\partial\chi}=k_{0}\chi\left(\sigma^{2}+\zeta^{2}+\delta^{2}\right)-k_{3}\left(\sigma^{2}-\delta^{2}\right)\zeta+\frac{2\chi}{\chi_{0}^{2}}\left[m_{\pi}^{2}f_{\pi}\sigma+\left(\sqrt{2}m_{K}^{2}f_{K}\,-\,\frac{1}{\sqrt{2}}m_{\pi}^{2}f_{\pi}\right)\zeta\right]$
$-\frac{\chi}{\chi_{0}^{2}}\left(m_{\omega}^{2}\omega^{2}+m_{\rho}^{2}\rho^{2}\right)+\left(4k_{4}+1+{\rm ln}\frac{\chi^{4}}{\chi_{0}^{4}}-\frac{4\xi}{3}{\rm ln}\left(\left(\frac{\left(\sigma^{2}-\delta^{2}\right)\zeta}{\sigma_{0}^{2}\zeta_{0}}\right)\right)\right)\chi^{3}=0,$ (55)
[…]
$-\,\sum_{i}g_{\delta i}\rho_{i}^{s}=0,~{}~{}~{}~{}~{}$ (56)
$\frac{\partial\Omega}{\partial\omega}=\left(\frac{\chi^{2}}{\chi_{0}^{2}}\right)m_{\omega}^{2}\omega+4g_{4}\omega^{3}+12g_{4}\omega\rho^{2}-\sum_{i}g_{\omega i}\rho_{i}=0,$ (57)
$\frac{\partial\Omega}{\partial\rho}=\left(\frac{\chi^{2}}{\chi_{0}^{2}}\right)m_{\rho}^{2}\rho+4g_{4}\rho^{3}+12g_{4}\omega^{2}\rho-\sum_{i}{g_{\rho i}}\rho_{i}=0,$ (58)
and
$\frac{\partial\Omega}{\partial\phi}=\left(\frac{\chi^{2}}{\chi_{0}^{2}}\right)m_{\phi}^{2}\phi+8g_{4}\phi^{3}-\sum_{i}g_{\phi\,i}\rho_{i}=0.$ (59)
In eqs. 57, 58 and 59 $\rho_{i}$ represents the vector density of baryons and is given by,
$\rho_{i}=\gamma_{i}\int\frac{d^{3}k}{(2\pi)^{3}}\Big{(}f_{i}(k)-\bar{f}_{i}(k)\Big{)},$ (60)
and in sections VI, VI and VI $\rho_{i}^{s}$ stands for the scalar density of baryons, given as
$\rho_{i}^{s}=\gamma_{i}\int\frac{d^{3}k}{(2\pi)^{3}}\frac{M_{i}^{*}}{E^{\ast}_{i}(k)}\Big{(}f_{i}(k)+\bar{f}_{i}(k)\Big{)}.$ (61)
Here, $f_{i}(k)$ and $\bar{f}_{i}(k)$ appearing in eqs. 60 and 61 represent the Fermi distribution functions for fermions and anti-fermions, at finite temperature and are expressed as
$f_{i}(k)$ $=$ $\frac{1}{1+\exp\left[(E^{*}_{i}(k)-\nu^{*}_{i})/k_{B}T\right]}~{}~{}~{}\text{and}~{}~{}~{}\bar{f}_{i}(k)=\frac{1}{1+\exp\left[(E^{*}_{i}(k)+\nu^{*}_{i})/k_{B}T\right]}~{}.$ (62)

### v2

@@ 3.2 In-medium magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances
A study based on QMC as well as MQMC, for octet baryons, shows that we can expect sizable influence on effective baryonic moments due to temperature above $T=150$ MeV at lower medium densities mqmc . Speaking of correlations, the authors in ref. seventyfour have discussed a strong correlation of the baryonic magnetic moment with that of the bag radius of a baryon, as the two quantities have a similar behavior as density dependent functions. Since this is the first work on medium modifications of $\frac{1}{2}^{-}$ baryon resonances, so referring to previous works on in-medium density effects on magnetic moments of baryons using the current formalism sixtytwo ; fiftyfour and QMC tsushima2 , might facilitate the purpose of a comparative study for future work on the baryonic magnetic moments of $\frac{1}{2}^{-}$ octet resonances. The effective masses of positive parity octet baryons are also observed to decrease as a function of nuclear density. As an example, the effective masses of proton $p$ is observed to decrease by 44 % as baryon density is decreased from 0 to $3\rho_{0}$ and this can be compared with the decrease of 25.27% for negative parity $p^{*}$ baryon. At high baryon density chiral symmetry is expected to restore partially and in this limit, properties of positive and negative parity baryons may converge to same values. This kind of picture can be seen more clearly within the present model calculations, if negative parity baryons are considered along with the positive parity baryons in the solution of system of non-linear equations corresponding to scalar and vector densities of baryons. As we discussed earlier, in the present work properties of negative parity baryons are investigated in strange matter comprising of positive parity nucleons and hyperons. In our future work, we shall extend our present model calculations to study the thermodynamics of dense matter considering explicitly both positive and negative baryons in equations of motion of scalar and vector fields through the definitions of scalar and vector densities of baryons.
4 Summary
@@
In a nutshell, this work investigates the medium modified trends induced on the masses and moreover on the magnetic moments of $J_{p}={\frac{1}{2}}^{-}$ baryon resonances in isospin asymmetric strange hadronic matter. The effects caused due to the presence of a medium are introduced through the CQMF model, following which we computed the effective baryon masses by using the masses of constituent quarks. The mass adjusted quark magnetic moments are then calculated, to include relativistic quark confinement, by using the baryon masses as inputs. In hadronic medium the in-medium masses of lights quarks, $u$ $d$, vary significantly in comparison to the observed changes in medium modified masses of $s$ quark. In the $\chi$CQM formalism, we then obtained the baryonic magnetic moments by the summation over the contributed values coming from valence quarks, sea quarks and orbital angular momentum due of sea quarks. As derived from the results is that, the medium modifications are more prominent for the baryons comprised of light $u$ or $d$ quarks, than for hyperons. For strange matter, the in-medium masses of constituent quark and baryons as well as the effective baryonic magnetic moments suffer appreciable enhancement in their in-medium behaviour. By carrying out this study at finite temperatures and finite $i_{a}$ values of the hadronic medium, we have noted a fine shift in the in-medium baryonic magnetic moments, although the temperature effects have been negligible below $T=150$ MeV.
The results from this study could be a point of reference to experiments which study hadron spectroscopy by probing photoproductions. In Jefferson Lab the GlueX experiment has measured the width of the N(1535) state to be narrower (less than 100 MeV) making it undetetctable, whereas its cascade program has aligned its goals to explore the negative parity excited $\Xi$ baryonsbaryon_spectroscopy . In the limelight of excited cascade resonances, a report from the Belle Collaboration sheds light on the detection of two states in low mass regime of which one ought to be $\Xi$(1690) Belle . Such experimental studies have research interests directed towards a detailed understanding of QCD for low-lying hadronic structures. Hence, the density dependent study of negative parity first excited states of octet baryon resonances is a crucial point of experimental interest at exploring the predictions of constituent quark models. Additionally, this work has scope for further extrapolation in the context of higher density astrophysical objects like neutron stars, to understand the impact of a dense medium on the baryonic magnetic moments medium_effect_neutron .
$\begin{array}[]{|c|c|c|c|c|c|c|}\hline\cr k_{0}&k_{1}&k_{2}&k_{3}&k_{4}\\ \hline\cr 4.94&2.12&-10.16&-5.38&-0.06\\ \hline\cr\sigma_{0}\,(\text{MeV})&\quad\zeta_{0}\,(\text{MeV})&\quad\chi_{0}\,(\text{MeV})&\quad\xi&\quad\rho_{0}\,(\text{fm}^{-3})\\ \hline\cr-92.8&-96.5&254.6&\frac{6}{33}&0.16\\ \hline\cr g_{\sigma}^{u}=g_{\sigma}^{d}&\quad g_{\sigma}^{s}=g_{\zeta}^{u}=g_{\zeta}^{d}&\quad g_{\delta}^{u}&g_{\zeta}^{s}=g_{s}&g_{4}\\ \hline\cr 2.72&0&2.72&3.847&37.4\\ \hline\cr g_{\rho}^{p}=g_{\rho}^{u}&g_{\omega}^{N}=3g_{\omega}^{u}&\quad g_{\rho}^{p}&\quad m_{\pi}\,(\text{MeV})&\quad m_{K}\,(\text{MeV})\\ \hline\cr 2.72&9.69&8.886&139&494\\ \hline\cr\end{array}$
Table 1: In above table various parameters used in the current manuscript are summarized.
@@
Table 4: Computed values of effective magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of symmetric nuclear matter ($i_{a}=0$ $f_{s}=0$) at $T=0$ MeV.
@@
${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.151 -0.367 0.196 1.980 2.477 -0.421 0.019 2.074
${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -1.463 0.263 -0.183 -1.382 -1.686 0.305 0.004 -1.376
@@
Table 5: Computed values of effective magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of asymmetric nuclear matter ($i_{a}=0.5$ $f_{s}=0$) at $T=0$ MeV.
@@
Table 6: Computed values of effective magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of symmetric hyperonic matter ($i_{a}=0$ $f_{s}=0.3$) at $T=0$ MeV.
@@
Table 7: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of asymmetric hyperonic matter ($i_{a}=0.5$ $f_{s}=0.3$) at $T=0$ MeV.
@@
$\mu_{\text{val}}$ $\mu_{\text{sea}}$ $\mu_{\text{orbit}}$ $\mu_{B}$ $\mu_{B}$ (NGBE) $\mu_{B}$ (OGE) $\mu_{B}$ (NGBE) $\mu_{B}$ $\mu_{B}$ $\mu_{B}$